What is the x and y intercept for y=2x + 5
Answer by stanbon(75887)

What is the x and y intercept for y=2x + 5
---
x-int ?
Let y = 0,
then 2x+5 = 0
x = -5/2
That is the x-intercept
---------------
y-int ?
Let x = 0
then y = 2*0+5
y = 5
That is the y-intercept
----------
Cheers,
Stan H.
--------